# Recursively-Constrained Partially Observable Markov Decision Processes

### v2

@@ Appendix B Proof of Lemma 1
Given a maximum one-step cost $C_{max}$ for $\pi_{c_{min}}$ and a non-negative admissible cost $v.d$, a lower bound on the admissible horizon can be obtained as follows. The $k$-step admissible horizon from each leaf node $v$ is the largest $k$ such that
@@
Also, we obtain the $\infty$-admissibility condition on $\pi_{c_{min}}$ by setting $k=\infty$ in Eq. (24):
@@
Next, when $v.\overline{V}_{C}{\pi_{c_{min}}}=0\leq v.d$, since costs are non-negative, the minimum cost policy obtains 00 cost at every future belief, and hence $k=\infty$.
@@
Appendix C Proof of Proposition 3
Proposition 3.
PRUNE only removes sub-optimal policies.
@@ Appendix C Proof of Proposition 3
Appendix D Proof of Lemma 2
Lemma 2 (Bound Validity).
Given an RC-POMDP with admissibility constraint $k=\infty$, let $T$ be the computed policy tree after some iterations of ARCS. Let $V_{R}^{*}$ and $V_{C}^{*}$ be the reward-value and cost-value of an optimal admissible policy. At every node $v$ with $\bar{b}=(v.b,v.d)$ with admissible horizon guarantee $v.k=\infty$, it holds that:
$v.\underline{V}_{R}\leq V_{R}^{*}(\bar{b})\leq v.\overline{V}_{R}\text{\;\; and \;\;}v.\underline{V}_{C}\leq V_{C}^{*}(\bar{b})\leq v.\overline{V}_{C}.$
@@ Appendix D Proof of Lemma 2
$\overline{V}_{R}$ and $\underline{V}_{C}$ are computed from the sawtooth approximation using $V^{U}_{R}$ and $V^{L}_{C}$, initialized with the Fast Informed Bound with the unconstrained POMDP problem, separately for reward maximization and cost minimization. The sawtooth approximation with the Fast Informed Bound provides valid upper bounds on reward-value (and lower bounds on cost-value of a cost-minimization policy, which in turn is a lower bound on cost-value for an optimal RC-POMDP policy) (Hauskrecht 2000). The upper bound on the optimal reward-value for the reward-maximization unconstrained POMDP problem is also an upper bound on the optimal reward-value for an RC-POMDP which has additional constraints. Similarly, the lower bound on the optimal cost-value for the cost-minimization unconstrained POMDP problem is also an lower bound on the optimal cost-value for an RC-POMDP which has additional constraints.
$\underline{V}_{R}$ and $\overline{V}_{C}$ are computed using the minimum cost policy $\Gamma_{c_{min}}$. This minimum cost policy is an alpha-vector policy, which is an upper bound on the cost-value function (Hauskrecht 2000), so $\overline{V}_{C}$ is an upper bound on the cost-value when following the minimum cost policy. It can also be seen that the value $\underline{V}_{R}$ from following the same policy is a valid bound on the optimal reward-value function from that node. Finally, the admissible horizon guarantee $k$ is initialized using the results from Lemma 1, which is shown to be a lower bound on the true admissible horizon following the minimum cost policy.
Next, we show that performing the BACKUP step (Algorithm 4 maintains the validity of the bounds. Recall the condition of this lemma that admissible horizon guarantee $v.k=\infty$ for the node $v$. Thus, after the backup step, the admissible horizon guarantee remains at $v.k=\infty$. From the proof of Theorem 3, the RC-POMDP Bellman backup $\mathbb{B}$ satisfies Bellman’s Principle of Optimality and is a contraction mapping within the space of admissible policies. Let $\overline{V}_{R}^{\prime},\underline{V}_{R}^{\prime},\overline{V}_{C}^{\prime},\underline{V}_{C}^{\prime}$ be the value after the BACKUP step, which performs a Bellman backup $\mathbb{B}$ on $\overline{V}_{R},\underline{V}_{R},\overline{V}_{C},\underline{V}_{C}$:
@@
Since $\mathbb{B}$ is a contraction mapping, we see that:
@@
$v.\underline{V}_{C}\leq V_{C}^{*}(\bar{b})\leq v.\overline{V}_{C}$
Appendix E Proof of Theorem 4
Theorem 4 (Soundness).
Given an RC-POMDP with admissibility constraint $k=\infty$ and $\epsilon>0$, if ARCS terminates with a solution, the solution is both an admissible policy and an $\epsilon$-optimal policy.
@@ Appendix E Proof of Theorem 4
Appendix F Experimental Evaluation
F.1 Implementation details
The code for each algorithm implementation can be found in the code appendix. Here, we detail parameters and implementation of the algorithms. For hyper-parameter tuning, we used the default parameters for ARCS. For the rest of the algorithms, the values of the hyper-parameters were chosen based empirical evaluations, and fixed for the experiments. For each environment, we used a maximum time step of $20$ during evaluation. Except for the Tiger problem, all algorithms reached terminal states before $20$ time steps in these problems or produced a policy which stayed still at $20$ time steps.
ARCS
We implemented ARCS as described. We used the Fast Informed Bound for the initialization of upper bound on reward value and lower bound on cost value. We used SARSOP for the computation of the minimum cost policy. We set the SARSOP hyperparameter $\kappa=0.5$ for our experiments, the same value as (Kurniawati, Hsu, and Lee 2008). We used a uniform randomization ($0.5$ probability) between heuristic sampling and random sampling during planning, and we leave an analysis on how the randomization weight may affect planning efficiency to future work.
CGCP
We implemented CGCP and adapted it for discounted infinite horizon problems, using Alg. 5 in (Walraven and Spaan 2018) as a basis. However, we use the discounted infinite horizon POMDP solver SARSOP (Kurniawati, Hsu, and Lee 2008) in place of a finite horizon PBVI. Our method of constructing policy graphs also differs, as the approach described is for finite horizon problems. We check for a return to beliefs previously visited under the policy in order to reduce the size of the graph. A maximum time of $300$ seconds was used for CGCP. For each SARSOP iteration within CGCP, $\tau=20$ seconds was given initially, while the solve time was incremented by $\tau^{+}=100$ seconds every time that the dual price $\lambda$ remained the same. Additionally, CGCP was limited to $100$ iterations. In an effort to reduce computation time, policy graph evaluation and SARSOP search was limited to depth $20$ (the same as the monte carlo evaluation depth) in all domains except the RockSample domains, which were allowed unlimited depth.
For the Tunnels benchmark, $1000$ monte carlo simulations to depth $20$ (the same as the evaluation depth) were used in place of a policy graph to estimate the value of policies. This was due to the inability of the policy graphs to estimate the value of some infinite horizon POMDP solutions which do not lead to terminal states or beliefs which have already appeared in the tree.
CGCP-CL
@@ F.1 Implementation details
CPBVI
We implemented CPBVI based on (Kim et al. 2011). The algorithm generates a set of reachable beliefs $\mathcal{B}$ before performing iterations of approximate dynamic programming on the belief set. However, the paper did not include full details on belief set $\mathcal{B}$ generation and alpha-vector set $\Gamma$ initialization.
The paper cited (Pineau et al. 2006) for their belief set description, and so we followed (Pineau et al., 2006) [^2] J. Pineau, G. Gordon, and S. Thrun. Anytime point-based approximations for large POMDPs. Journal of Artificial Intelligence Research, 27:335–380, 2006. by expanding $\mathcal{B}$ greedily towards achieving uniform density in the set of reachable beliefs. This is done by randomly simulating a step forward from a node in the tree, thereby generating candidate beliefs, and keeping the belief that is farthest from any belief already in the tree. We repeat this expansion until the desired number of beliefs have been added to the tree.
@@
The CPBVI algorithm involves the computation of a linear program (LP) to obtain the best action at a given belief. One of the constraints asserts that the convex combination of cost alpha-vectors evaluated at a given belief $b$ must be equal to or less than the admissible cost $d$ associated with $b$, which is used in CPBVI’s heuristic approach. However, if $d<0$, the LP becomes infeasible. The case of $d<0$ is possible since no pruning of beliefs is conducted. The paper did not provide details to account for this situation. To address this, if the LP is infeasible, we output the action with the lowest cost, akin to ARC’s minimum cost policy method when no policy is admissible.
CPBVI-D
@@
F.2 Environment details
@@
Tunnels Problem
The tunnels problem is modeled as an RC-POMDP as follows. As depicted in figure 2, the tunnels problem consists of a centralized starting hall that funnels into 3 separate tunnels. At the end of tunnels 1,2,and 3 lie rewards 2.0, 1.5 and 0.5 respectively. However, with high reward also comes high cost as tunnel 1 has a 60% probability of containing rocks and tunnel 2 has a 40% probability of containing rocks while tunnel 3 is always free of rocks. If present, the rocks fill 2 steps before the reward location at the end of a tunnel and a cost of 1 is incurred if the agent traverses over these rocks. Furthermore, a cost of 1 is incurred if the agent chooses to move backwards.
@@
F.3 Experiment Results
We implemented each algorithm in Julia using tools from the POMDPs.jl framework (Egorov et al. 2017), and all experiments were conducted on a computer with two nominally 2.2 GHz Intel Xeon CPUs with 48 cores and 128 GB RAM. All experiments were conducted in Ubuntu 18.04.6 LTS. The full results with the mean and standard error of the mean for each metric are shown in Table. 2.
Environment Algorithm Violation Rate Cumulative Reward Cumulative Cost
@@ F.3 Experiment Results
Table 2: Comparison of our RC-POMDP algorithm to state-of-the-art offline C-POMDP algorithms. We report the mean and $1$ standard error of the mean for each metric.

### v3

@@ Appendix D Proof of Lemma 1
Given a maximum one-step cost $C_{max}$ for $\pi_{c}^{min}$ and a non-negative admissible cost $v.d$, a lower bound on the admissible horizon can be obtained as follows. The $k$-step admissible horizon from each leaf node $v$ is the largest $k$ such that
@@
Also, we obtain the $\infty$-admissibility condition on $\pi_{c}^{min}$ by setting $k=\infty$ in Eq. (23):
@@
Next, when $v.\overline{V}_{C}^{\pi_{c^{min}}}=0\leq v.d$, since costs are non-negative, the minimum cost policy obtains 00 cost at every future belief, and hence $k=\infty$.
@@
Appendix E Proof of Proposition 3
@@
Appendix F Proof of Lemma 2
@@
$\overline{V}_{R}$ and $\underline{V}_{C}$ are initialized with the Fast Informed Bound with the unconstrained POMDP problem, separately for reward maximization and cost minimization. The Fast Informed Bound provides valid upper bounds on reward-value (and lower bounds on cost-value of a cost-minimization policy, which in turn is a lower bound on cost-value for an optimal RC-POMDP policy) Hauskrecht [2000]. The upper bound on the optimal reward-value for the reward-maximization unconstrained POMDP problem is also an upper bound on the optimal reward-value for an RC-POMDP which has additional constraints. Similarly, the lower bound on the optimal cost-value for the cost-minimization unconstrained POMDP problem is also a lower bound on the cost-value of an optimal RC-POMDP policy which has additional constraints.
$\underline{V}_{R}$ and $\overline{V}_{C}$ are computed using the minimum cost policy $\Gamma_{c_{min}}$. This minimum cost policy is an alpha-vector policy, which is an upper bound on the cost-value function Hauskrecht [2000], so $\overline{V}_{C}$ is an upper bound on the cost-value when following the minimum cost policy. It can also be seen that the value $\underline{V}_{R}$ from following the same policy is a valid bound on the optimal reward-value function from that node. Finally, the admissible horizon guarantee $k$ is initialized using the results from Lemma 1, which is shown to be a lower bound on the true admissible horizon following the minimum cost policy.
Next, we show that performing the BACKUP step (Algorithm 4 maintains the validity of the bounds. Recall the condition of this lemma that admissible horizon guarantee $v.k=\infty$ for the node $v$. Thus, after the backup step, the admissible horizon guarantee remains at $v.k=\infty$. From the proof of Theorem 2, the RC-POMDP Bellman backup $\mathbb{B}$ satisfies Bellman’s Principle of Optimality and is a contraction mapping within the space of admissible value functions (and hence policies).
Let $\overline{V}_{R}^{\prime},\underline{V}_{R}^{\prime},\overline{V}_{C}^{\prime},\underline{V}_{C}^{\prime}$ be the value after the BACKUP step, which performs a Bellman backup $\mathbb{B}$ on $\overline{V}_{R},\underline{V}_{R},\overline{V}_{C},\underline{V}_{C}$:
@@
Since $\mathbb{B}$ is a contraction mapping within the space of admissible policies, we see that:
@@
$V_{C}^{*}(\bar{b})\leq v.\overline{V}_{C}$
Appendix G Proof of Theorem 4
@@
Appendix H Experimental Evaluation
H.1 Implementation details
The code for each algorithm implementation can be found in the attached supplementary material. Here, we detail parameters and implementation of the algorithms. For hyper-parameter tuning, we used the default parameters for ARCS. For the rest of the algorithms, the values of the hyper-parameters were chosen based empirical evaluations, and fixed for the experiments. For each environment, we used a maximum time step of $20$ during evaluation. Except for the Tiger problem, all algorithms reached terminal states before $20$ time steps in these problems or produced a policy which stayed still at $20$ time steps.
H.1.1 ARCS
We implemented ARCS as described. We used the Fast Informed Bound for the initialization of upper bound on reward value and lower bound on cost value. We used SARSOP for the computation of the minimum cost policy. We set the SARSOP hyperparameter $\kappa=0.5$ for our experiments, the same value as Kurniawati et al. [2008]. We used a uniform randomization ($0.5$ probability) between heuristic sampling and random sampling during planning, and we leave an analysis on how the randomization weight may affect planning efficiency to future work.
H.1.2 CGCP
We implemented CGCP and adapted it for discounted infinite horizon problems, using Alg. 5 in Walraven and Spaan [2018] as a basis. However, we use the discounted infinite horizon POMDP solver SARSOP Kurniawati et al. [2008] in place of a finite horizon PBVI. Our method of constructing policy graphs also differs, as the approach described is for finite horizon problems. We check for a return to beliefs previously visited under the policy in order to reduce the size of the graph. A maximum time of $300$ seconds was used for CGCP. For each SARSOP iteration within CGCP, $\tau=20$ seconds was given initially, while the solve time was incremented by $\tau^{+}=100$ seconds every time that the dual price $\lambda$ remained the same. Additionally, CGCP was limited to $100$ iterations. In an effort to reduce computation time, policy graph evaluation and SARSOP search was limited to depth $20$ (the same as the monte carlo evaluation depth) in all domains except the RockSample domains, which were allowed unlimited depth.
For the Tunnels benchmark, $1000$ monte carlo simulations to depth $20$ were used in place of a policy graph to estimate the value of policies. This was due to the inability of the policy graphs to estimate the value of some infinite horizon POMDP solutions which do not lead to terminal states or beliefs which have already appeared in the tree.
H.1.3 CGCP-CL
@@
H.1.4 No-regret Learning Algorithm
We implemented the no-regret learning algorithm from [Kalagarla et al., 2022]. We used SARSOP as the unconstrained POMDP solver, and monte carlo simulations to estimate the value of policies.
H.1.5 CPBVI
We implemented CPBVI based on Kim et al. [2011]. The algorithm generates a set of reachable beliefs $\mathcal{B}$ before performing iterations of approximate dynamic programming on the belief set. However, the paper did not include full details on belief set $\mathcal{B}$ generation and alpha-vector set $\Gamma$ initialization.
The paper cited Pineau et al. [2006] for their belief set description, and so we followed Pineau et al. [2006] by expanding $\mathcal{B}$ greedily towards achieving uniform density in the set of reachable beliefs. This is done by randomly simulating a step forward from a node in the tree, thereby generating candidate beliefs, and keeping the belief that is farthest from any belief already in the tree. We repeat this expansion until the desired number of beliefs have been added to the tree.
@@ H.1.5 CPBVI
The CPBVI algorithm involves the computation of a linear program (LP) to obtain the best action at a given belief. One of the constraints asserts that the convex combination of cost alpha-vectors evaluated at a given belief $b$ must be equal to or less than the admissible cost $d$ associated with $b$, which is used in CPBVI’s heuristic approach. However, if $d<0$, the LP becomes infeasible. The case of $d<0$ is possible since no pruning of beliefs is conducted. The paper did not provide details to account for this situation. To address this, if the LP is infeasible, we output the action with the lowest cost, akin to ARCS’ minimum cost policy method when no policy is admissible.
H.1.6 CPBVI-D
@@
H.2 Environment details
  • CE: Simplified counterexample in Figure 1.
  • C-Tiger: A constrained version of the Tiger POMDP Problem Kaelbling et al. [1998], with cost of $1$ for the “listen" action.
  • CRS: A constrained version of the RockSample problem Smith and Simmons [2004] as defined in Lee et al. [2018] with varying sizes and number of rocks.
  • Tunnels: A scaled version of Example 1, shown in Fig. 2.
@@ H.2.1 Counterexample Problem
H.2.2 Tunnels Problem
The tunnels problem is modeled as an RC-POMDP as follows. As depicted in figure 2, the tunnels problem consists of a centralized starting hall that funnels into 3 separate tunnels. At the end of tunnels 1,2,and 3 lie rewards 2.0, 1.5 and 0.5 respectively. However, with high reward also comes high cost as tunnel 1 has a 80% probability of containing rocks and tunnel 2 has a 40% probability of containing rocks while tunnel 3 is always free of rocks. If present, the rocks fill 2 steps before the reward location at the end of a tunnel and a cost of 1 is incurred if the agent traverses over these rocks. Furthermore, a cost of 1 is incurred if the agent chooses to move backwards.
@@ H.2.2 Tunnels Problem
H.3 Experiment Evaluation Setup
We implemented each algorithm in Julia using tools from the POMDPs.jl framework Egorov et al. [2017], and all experiments were conducted single-threaded on a computer with two nominally 2.2 GHz Intel Xeon CPUs with 48 cores and 128 GB RAM. All experiments were conducted in Ubuntu 18.04.6 LTS. For all algorithms except CGCP-CL, solve time is limited to $300$ seconds and online action selection to $0.05$ seconds. For CGCP-CL, $300$ seconds was given for each action (recomputed from scratch). We simulate each policy $1000$ times, except for CGCP-CL, which is simulated $100$ times due to the time taken for re-computation of the policy at each time step. The full results with the mean and standard error of the mean for each metric are shown in Table. 3.
Table 3: Comparison of our RC-POMDP algorithm to state-of-the-art offline C-POMDP algorithms. The second column shows the size of the states, actions, and observations of each environment. We report the mean and $1$ standard error of the mean for each metric. A memory-out is indicated by a --. Note that for CRS(11,11), due to the $300s$ time limit, CGCP, Exp-Gradient and ours all compute a policy that goes directly to the exit area without interacting with any rocks, and achieve the same reward and 00 cost.
Environment State/Action/Obs Algorithm Violation Rate Cumulative Reward Cumulative Cost
@@ H.3 Experiment Evaluation Setup
CRS(7,8) CGCP $0.357\pm 0.015$ $10.78\pm 0.19$ $0.945\pm 0.04$
CL-CGCP $0.20\pm 0.13$ $11.17\pm 1.53$ $0.931\pm 0.078$
($\hat{c}=1$) 12545 / 13 / 3 CPBVI $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
CPBVI-D $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
EXP-Gradient $0.322\pm 0.015$ $10.03\pm 0.16$ $1.154\pm 0.054$
Ours $0.0\pm 0$ 0.0 $6.61\pm 0.22$ $0.960\pm 0.003$
CRS(11,11) CGCP $0.0\pm 0$ 0.0 $5.987\pm 0.0$ $0.0\pm 0$ 0.0
CL-CGCP - - -
($\hat{c}=1$) 247809 / 16 / 3 CPBVI $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
CPBVI-D $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
EXP-Gradient $0.0\pm 0$ 0.0 $5.987\pm 0.0$ $0.0\pm 0$ 0.0
Ours $0.0\pm 0$ 0.0 $5.987\pm 0.0$ $0.0\pm 0$ 0.0